# Label propagation on binomial random graphs

### v4

@@ 1 Introduction
Coincidentally, several recent articles have studied the binary case of majority dynamics when the underlying graph is $\mathcal{G}(n,p)$ and initially every vertex chooses one of two labels with equal probability (in contrast to labels from an interval as in LPA). The question considered is whether all vertices converge to the same label and, if so, how many rounds it takes. Benjamini, Chan, O’Donnell, Tamuz and Tan [3] showed that if $p=\Omega(1/\sqrt{n})$, then $\mathcal{G}(n,p)$ is such that with probability at least $0.4$ (over the choice of the random graph and the initial choice of vertex labels), convergence to the most popular label happens after four rounds. In fact, they conjectured that this holds with high probability. The conjecture was recently positively resolved by Fountoulakis, Kang and Makai [11] (see also [27, 32, 30]). Also, see [5] for recent work on majority dynamics in $\mathcal{G}(n,p)$ in the case where vertices are initially assigned one of $k=O(1)$ labels uniformly at random.
@@ 2.2 Preliminaries
and define $p_{Y}(M,\rho)$ similarly by exchanging $X$ and $Y$. Then, for every integer $M\geq 0$,
@@ 3.3 The regime $n^{5/8+\varepsilon}\leq np\ll n^{2/3}$
In this section, we present the delayed proof of Lemma 2.6. First of all, by rewriting the claimed inequality in the equivalent form
@@
Using that $p_{X}(M-1,\rho)\geq p_{X}(M,\rho)$ and $p_{Y}(M-1,\rho)\geq p_{Y}(M,\rho)$ for all integer $M\geq 0$ and real $\rho\geq 0$, (56) is equivalent to
@@
To begin with, (58) is clear since $p_{Y}(n,\rho)=0$ while all other terms are positive real numbers. Set $N=M+i-1$. Then, the left-hand side of (59) rewrites as
@@
$A(X)=$ $2\binom{n}{N}\binom{n-1}{N}X^{2N}+\rho\binom{n}{N+1}\binom{n-1}{N+1}X^{2N+2},$
$B(X)=$ $2\binom{n}{N-1}\binom{n-1}{N-1}X^{2N-2}+\rho\binom{n}{N}\binom{n-1}{N}X^{2N},$
@@
Note that actually $P_{1}=R_{2}$ and $Q_{1}=S_{2}$, but we nevertheless keep both notations for simplicity of the exposition.
To show that $P(X)S(X)\geq Q(X)R(X)$ (which is equivalent to (59)), we will consecutively show the following simpler inequalities for every $X\geq 0$:
@@
    $P_{1}(X)S_{2}(X)+P_{2}(X)S_{1}(X)\geq Q_{1}(X)R_{2}(X)+Q_{2}(X)R_{1}(X)$,
  4. I4
    $(P_{1}(X)-Q_{1}(X))B(X)\geq(R_{1}(X)-S_{1}(X))A(X)$,
  5. I5
    $(P_{2}(X)-Q_{2}(X))B(X)\geq(R_{2}(X)-S_{2}(X))A(X)$.
Step 1: showing I1.
We will show that all coefficients of the polynomial $P_{1}(X)S_{1}(X)-Q_{1}(X)R_{1}(X)$ are positive. First, we take care of the terms $X^{2N+k}$ with $k\in[0,N-2]$. For any such $k$, the coefficient in front of $X^{2N+k}$ rewrites
@@
Next, we turn to the coefficients in front of the terms $X^{2N+k}$ for $k\in[N-1,2N-3]$ where $i$ could be equal to $N-1$ but $j$ cannot, hence the loss of the symmetry used above. Nevertheless, using that the term corresponding to $i=N-1,j=k-(N-1)$ is the only term without symmetric counterpart, we obtain that the coefficient in front of $X^{2N+k}$ rewrites as a sum of two terms. Following (61), the first term contains the ‘symmetric part’ of the sum, which is equal to
@@
$\binom{n}{N}\binom{n}{N-1}\binom{n}{N-1}\binom{n}{k-N+1}\sum_{i,j\leq N-2,\,i+j=k}\frac{\tbinom{n}{i}\tbinom{n}{j}}{\tbinom{n}{N-1}\tbinom{n}{k-N+1}}\frac{n-k/2}{n^{2}}.$
@@
=\;= $\binom{n}{N}\binom{n}{N-1}^{2}\binom{n}{k-N+1}\frac{1-(n-N)(2N-k-3)}{n^{2}}.$
@@
which finishes the proof in this case.
Step 2: showing I2.
The proof of this second step follows from the proof of the first step by replacing $N$ with $N+1$ in all relevant places.
Step 3: showing I3.
First, since $P_{1}=R_{2}$ and $Q_{1}=R_{2}$, we have that $P_{1}(X)S_{2}(X)=Q_{1}(X)R_{2}(X)$. Thus, it remains to study the difference $P_{2}(X)S_{1}(X)-Q_{2}(X)R_{1}(X)$. As in step 1, we begin by studying the terms $X^{2N+k}$ with $k\in[0,N-2]$. For any such $k$, the coefficient in front of $X^{2N+k}$ rewrites
@@
Next, we turn to the coefficients in front of the terms $X^{2N+k}$ for $k\in[N-1,2N-2]$ where $i$ could be equal to $N-1$ or $N$ but $j$ cannot, hence the loss of the symmetry used above. We consider two subcases.
@@
$\sum_{i,j\geq 1,\,i+j=N-1}\binom{n}{N+1}\binom{n}{i}\binom{n}{N-1}\binom{n}{j}\frac{2n-k}{n^{2}},$
@@
=\;= $\binom{n}{N+1}\binom{n}{N-1}^{2}\frac{4-(n-N-1)(N-3)}{n^{2}}.$
@@
$\sum_{i,j\geq 1,\,i+j=N-1}\frac{2n-k}{n^{2}}=\frac{(N-2)(2n-N+1)}{n^{2}}\geq\frac{(N-3)(n-N-1)}{n^{2}}$
finishes the proof in this case.
Case 3.2: $k\in[N,2N-2]$.
@@
=\;= $\binom{n}{N+1}\binom{n}{N-1}^{2}\binom{n}{k-N+1}\frac{4-(n-N-1)(2N-4-k)}{n^{2}}.$ (62)
@@
=\;= $\binom{n}{N+1}\binom{n}{N}\binom{n}{N-1}\binom{n}{k-N}\frac{2-(n-N-1)(2N-2-k)}{n^{2}}.$ (63)
Now, suppose that $k\in\{2N-3,2N-2\}$. If $k=2N-2$, both  (62) and (63) are positive. Moreover, if $k=2N-3$, (62) is again positive and the inequality $\tbinom{n}{N-1}\tbinom{n}{N-2}\geq\tbinom{n}{N}\tbinom{n}{N-3}$ ensures that the sum of (62) and (63) remains positive.
Suppose that $k\in[N,2N-4]$. Again, minimizing the expression $\tbinom{n}{i}\tbinom{n}{j}$ under the constraints $i,j\leq N-2$ and $i+j=k$ requires $|i-j|$ to be largest possible (that is, $i=N-2$ and $j=k-N+2$ or vice versa). As a result, $\tbinom{n}{i}\tbinom{n}{j}\geq\tbinom{n}{N-1}\binom{n}{k-N+1}\geq\tbinom{n}{N}\binom{n}{k-N}$ for all $i,j\leq N-2$ with $i+j=k$. Combining this observation with the inequality
@@
Step 4: showing I4.
First, we simplify $(P_{1}(X)-Q_{1}(X))B(X)$, which rewrites as
$\bigg{(}\sum_{i=0}^{N-1}\bigg{(}\binom{n}{N}\binom{n-1}{i}$ $-\binom{n-1}{N}\binom{n}{i}\bigg{)}X^{N+i}\bigg{)}$
$\cdot\bigg{(}2\binom{n}{N-1}\binom{n-1}{N-1}X^{2N-2}+\rho\binom{n}{N}\binom{n-1}{N}X^{2N}\bigg{)}.$
Using once again that, for every $\ell\geq 0$, $\tbinom{n-1}{\ell}=\tfrac{n-\ell}{n}\tbinom{n}{\ell}$, we obtain that the above polynomial is equal to
$\begin{split}\bigg{(}\sum_{i=0}^{N-1}\binom{n}{N}\binom{n}{i}\frac{N-i}{n}&\cdot 2\binom{n}{N-1}\binom{n-1}{N-1}X^{3N+i-2}\bigg{)}\\ &+\bigg{(}\sum_{i=0}^{N-1}\binom{n}{N}\binom{n}{i}\frac{N-i}{n}\cdot\rho\binom{n}{N}\binom{n-1}{N}X^{3N+i}\bigg{)}.\end{split}$ (65)
Now, we simplify $(R_{1}(X)-S_{1}(X))A(X)$. This polynomial rewrites as
$\bigg{(}\sum_{i=0}^{N-2}\bigg{(}\binom{n}{N-1}\binom{n-1}{i}$ $-\binom{n-1}{N-1}\binom{n}{i}\bigg{)}X^{N+i-1}\bigg{)}$
$\cdot\bigg{(}2\binom{n}{N}\binom{n-1}{N}X^{2N}+\rho\binom{n}{N+1}\binom{n-1}{N+1}X^{2N+2}\bigg{)}.$
Using once again that, for every $\ell\geq 0$, $\tbinom{n-1}{\ell}=\tfrac{n-\ell}{n}\tbinom{n}{\ell}$, we obtain that the above polynomial rewrites
$\bigg{(}\sum_{i=0}^{N-2}\binom{n}{N-1}$ $\binom{n}{i}\frac{N-1-i}{n}\cdot 2\binom{n}{N}\binom{n-1}{N}X^{3N+i-1}\bigg{)}$
$+\bigg{(}\sum_{i=0}^{N-2}\binom{n}{N-1}\binom{n}{i}\frac{N-1-i}{n}\cdot\rho\binom{n}{N+1}\binom{n-1}{N+1}X^{3N+i+1}\bigg{)}.$
Furthermore, after change of variables, the latter expression further rewrites
$\begin{split}\bigg{(}\sum_{i=1}^{N-1}\binom{n}{N-1}&\binom{n}{i-1}\frac{N-i}{n}\cdot 2\binom{n}{N}\binom{n-1}{N}X^{3N+i-2}\bigg{)}\\ &+\bigg{(}\sum_{i=1}^{N-1}\binom{n}{N-1}\binom{n}{i-1}\frac{N-i}{n}\cdot\rho\binom{n}{N+1}\binom{n-1}{N+1}X^{3N+i}\bigg{)}.\end{split}$ (66)
Finally, we show that the coefficients of (65) dominate the coefficients of (66). Observe that, for every $i\in[N-1]$,
$\binom{n}{N}\binom{n}{i}\binom{n}{N-1}\binom{n-1}{N-1}-\binom{n}{N-1}\binom{n}{i-1}\binom{n}{N}\binom{n-1}{N}$
can be simplified to
$\binom{n}{N}\binom{n}{N-1}\binom{n}{i}\binom{n}{N}\bigg{(}\frac{N}{n}-\frac{i(n-N)}{(n-i+1)n}\bigg{)}\geq 0,$
so the first term in (65) dominates the first term in (66). Also, for every $i\in[N-1]$,
$\binom{n}{N}\binom{n}{i}\binom{n}{N}\binom{n-1}{N}-\binom{n}{N-1}\binom{n}{i-1}\binom{n}{N+1}\binom{n-1}{N+1}$
simplifies to
$\binom{n}{N}\binom{n}{i}\binom{n}{N+1}\binom{n-1}{N+1}\bigg{(}\frac{(N+1)^{2}}{(n-N)(n-N-1)}-\frac{Ni}{(n-N+1)(n-i+1)}\bigg{)}.$
Using that $N+1\geq N\geq i$ and $n-i+1\geq n-N-1$ implies the positivity of the latter expression, which in turn means that the second term of (65) dominates the second term of (66)), thus confirming I4.
Step 5: showing I5.
@@
The inequality I5 follows similarly to the proof of the fourth step using that $P_{2},Q_{2},R_{2},S_{2}$ are obtained from $P_{1},Q_{1},R_{1},S_{1}$ by replacing $N$ with $N+1$.

### v5

@@ 1 Introduction
Coincidentally, several recent articles have studied the binary case of majority dynamics when the underlying graph is $\mathcal{G}(n,p)$ and initially every vertex chooses one of two labels with equal probability (in contrast to labels from an interval as in LPA). The question considered is whether all vertices converge to the same label and, if so, how many rounds it takes. Benjamini, Chan, O’Donnell, Tamuz and Tan [3] showed that if $p=\Omega(1/\sqrt{n})$, then $\mathcal{G}(n,p)$ is such that with probability at least $0.4$ (over the choice of the random graph and the initial choice of vertex labels), convergence to the most popular label happens after four rounds. In fact, they conjectured that this holds with high probability. The conjecture was recently positively resolved by Fountoulakis, Kang and Makai [11] (see also [28, 34, 31]). Also, see [6] for recent work on majority dynamics in
G(n,p)
[…]
in the case where vertices are initially assigned one of $k=O(1)$ labels uniformly at random.
@@ 2.2 Preliminaries
and define $p_{Y}(M,\rho)$ similarly by exchanging $X$ and $Y$. Then, for every integer $M\in[0,n]$,
@@ 3.3 The regime $n^{5/8+\varepsilon}\leq np\ll n^{2/3}$
In this proof, we work with an integer $M\in[0,n]$ and under the convention that, when a fraction has a positive numerator and denominator 00, it is equal to infinity. In this section, we present the delayed proof of Lemma 2.6. First of all, by rewriting the claimed inequality in the equivalent form
@@
Using that $p_{X}(M-1,\rho)>p_{X}(M,\rho)$ and $p_{Y}(M-1,\rho)>p_{Y}(M,\rho)$ for all integer $M\in[0,n-1]$ and real $\rho\geq 0$, (56) is equivalent to
@@
Again, (58) is clear since $p_{Y}(n,\rho)=0$ while all other terms are positive real numbers. Set $N=M+i-1$. Then, the left-hand side of (59) rewrites as
@@
$A(x)=A_{1}(x)+\rho A_{2}(x)\qquad\text{and}\qquad B(x)=B_{1}(x)+\rho B_{2}r(x)$
with
$A_{1}(x)$ $=2\binom{n}{N}\binom{n-1}{N}x^{2N},\quad$ $A_{2}(x)=\binom{n}{N+1}\binom{n-1}{N+1}x^{2N+2},$
$B_{1}(x)$ $=2\binom{n}{N-1}\binom{n-1}{N-1}x^{2N-2},\quad$ $B_{2}(x)=\binom{n}{N}\binom{n-1}{N}x^{2N},$
@@
Note that actually $P_{1}(x)=R_{2}(x)$ and $Q_{1}(x)=S_{2}(x)$, but we nevertheless keep both notations for simplicity of the exposition. In the sequel, we abbreviate notation and write $P_{1},Q_{1},\ldots$ instead of $P_{1}(x),Q_{1}(x),\ldots$, leaving the dependence on $x$ implicit for better readability.
To derive (59), we show that, for all $x,\rho\geq 0$, $PS\geq QR$ or, equivalently, that
$(\rho+1)^{2}(2\rho+4)PS\geq(\rho+1)^{2}(2\rho+4)QR.$ (60)
Observe that
$(\rho+1)^{2}(2\rho+4)PS$
$=P_{1}\big{[}(2\rho+4)S_{1}+\rho(2\rho+4)S_{2}+(\rho+1)B_{1}+\rho(\rho+1)B_{2}\big{]}$
$\qquad+P_{2}\big{[}\rho(2\rho+4)S_{1}+\rho^{2}(2\rho+4)S_{2}+\rho(\rho+1)B_{1}+\rho^{2}(\rho+1)B_{2}]$
$\qquad+(A_{1}+\rho A_{2})\big{[}(\rho+1)S_{1}+\rho(\rho+1)S_{2}\big{]}+(\rho+1)^{2}AB/(2\rho+4)$
$=\rho^{3}\cdot\big{[}2P_{2}S_{2}+P_{2}B_{2}+S_{2}A_{2}]$
$\qquad+\rho^{2}\cdot\big{[}2P_{1}S_{2}+P_{1}B_{2}+2P_{2}S_{1}+4P_{2}S_{2}+P_{2}B_{1}+P_{2}B_{2}+S_{1}A_{2}+S_{2}A_{1}+S_{2}A_{2}\big{]}$
$\qquad+\rho\cdot\big{[}2P_{1}S_{1}+4P_{1}S_{2}+P_{1}B_{1}+P_{1}B_{2}+4P_{2}S_{1}+P_{2}B_{1}+S_{1}A_{1}+S_{1}A_{2}+S_{2}A_{1}\big{]}$
$\qquad+1\cdot\big{[}4P_{1}S_{1}+P_{1}B_{1}+S_{1}A_{1}\big{]}+(\rho+1)^{2}AB/(2\rho+4),$
and similarly
$(\rho+1)^{2}(2\rho+4)QR$
$=\rho^{3}\cdot\big{[}2Q_{2}R_{2}+Q_{2}B_{2}+R_{2}A_{2}]$
$\qquad+\rho^{2}\cdot\big{[}2Q_{1}R_{2}+Q_{1}B_{2}+2Q_{2}R_{1}+4Q_{2}R_{2}+Q_{2}B_{1}+Q_{2}B_{2}+R_{1}A_{2}+R_{2}A_{1}+R_{2}A_{2}\big{]}$
$\qquad+\rho\cdot\big{[}2Q_{1}R_{1}+4Q_{1}R_{2}+Q_{1}B_{1}+Q_{1}B_{2}+4Q_{2}R_{1}+Q_{2}B_{1}+R_{1}A_{1}+R_{1}A_{2}+R_{2}A_{1}\big{]}$
$\qquad+1\cdot\big{[}4Q_{1}R_{1}+Q_{1}B_{1}+R_{1}A_{1}\big{]}+(\rho+1)^{2}AB/(2\rho+4),$
The remainder of this proof is dedicated to showing the following simpler auxiliary inequalities for every $x\geq 0$:
@@
    $P_{1}S_{2}+P_{2}S_{1}\geq Q_{1}R_{2}+Q_{2}R_{1}$,
and use them to derive that, for every $x\geq 0$, each of the following inequalities holds as well:
  1. J1
    $2P_{2}S_{2}+P_{2}B_{2}+S_{2}A_{2}\geq 2Q_{2}R_{2}+Q_{2}B_{2}+R_{2}A_{2}$,
  2. J2
    $2P_{1}S_{2}+P_{1}B_{2}+2P_{2}S_{1}+4P_{2}S_{2}+P_{2}B_{1}+P_{2}B_{2}+S_{1}A_{2}+S_{2}A_{1}+S_{2}A_{2}$
                                  $\geq 2Q_{1}R_{2}+Q_{1}B_{2}+2Q_{2}R_{1}+4Q_{2}R_{2}+Q_{2}B_{1}+Q_{2}B_{2}+R_{1}A_{2}+R_{2}A_{1}+R_{2}A_{2}$,
  3. J3
    $2P_{1}S_{1}+4P_{1}S_{2}+P_{1}B_{1}+P_{1}B_{2}+4P_{2}S_{1}+P_{2}B_{1}+S_{1}A_{1}+S_{1}A_{2}+S_{2}A_{1}$
                                  $\geq 2Q_{1}R_{1}+4Q_{1}R_{2}+Q_{1}B_{1}+Q_{1}B_{2}+4Q_{2}R_{1}+Q_{2}B_{1}+R_{1}A_{1}+R_{1}A_{2}+R_{2}A_{1}$,
  4. J4
    $4P_{1}S_{1}+P_{1}B_{1}+S_{1}A_{1}\geq 4Q_{1}R_{1}+Q_{1}B_{1}+R_{1}A_{1}$.
Showing I1.
We will show that all coefficients of the polynomial $P_{1}S_{1}-Q_{1}R_{1}$ are positive. First, we consider the terms $x^{2N-1+k}$ with $k\in[0,N-2]$. For any such $k$, the coefficient in front of $x^{2N-1+k}$ rewrites
@@
Next, we turn to the coefficients in front of the terms $x^{2N-1+k}$ for $k\in[N-1,2N-3]$ where $i$ could be equal to $N-1$ but $j$ cannot, hence the loss of the symmetry used above. Nevertheless, using that the term corresponding to $i=N-1,j=k-(N-1)$ is the only term without symmetric counterpart, we obtain that the coefficient in front of $x^{2N-1+k}$ rewrites as a sum of two terms. Following (62), the first term contains the ‘symmetric part’ of the sum, which is equal to
@@
$\binom{n}{N}\binom{n}{N-1}\binom{n}{N-1}\binom{n}{k-N+1}\sum_{i,j\leq N-2,\,i+j=k}\frac{\tbinom{n}{i}\tbinom{n}{j}}{\tbinom{n}{N-1}\tbinom{n}{k-N+1}}\frac{n-k/2}{n^{2}}.$ (63)
@@
$=\binom{n}{N}\binom{n}{N-1}^{2}\binom{n}{k-N+1}\frac{1-(n-N)(2N-k-3)}{n^{2}}.$ (64)
@@
which shows that the sum of (63) and (64) is positive and finishes the proof of I1.
Showing I2.
The proof of this second step follows from the proof of the first step by replacing $N$ with $N+1$ in all places.
Showing I3.
First, since $P_{1}=R_{2}$ and $Q_{1}=S_{2}$, we have that $P_{1}S_{2}=Q_{1}R_{2}$. Thus, it remains to study the difference $P_{2}S_{1}-Q_{2}R_{1}$. As for I1, we begin by studying the terms $x^{2N+k}$ with $k\in[0,N-2]$. For any such $k$, the coefficient in front of $x^{2N+k}$ rewrites as
@@
Next, we turn to the coefficients in front of the terms $x^{2N+k}$ for $k\in[N-1,2N-2]$ where $i$ could be equal to $N-1$ or $N$ but $j$ cannot, hence the loss of the symmetry used above. We consider three subcases.
@@
$\sum_{i,j\geq 1,\,i+j=N-1}\binom{n}{N+1}\binom{n}{i}\binom{n}{N-1}\binom{n}{j}\frac{2n-k}{n^{2}},$ (65)
@@
$=\binom{n}{N+1}\binom{n}{N-1}^{2}\frac{4-(n-N-1)(N-3)}{n^{2}}.$ (66)
@@
$\sum_{i,j\geq 1,\,i+j=N-1}\frac{2n-k}{n^{2}}=\frac{(N-2)(2n-N+1)}{n^{2}}\geq\frac{(N-3)(n-N-1)}{n^{2}},$
showing that the sum of (65) and (66) is positive and finishing the proof in this case.
Case 3.2: $k\in[N,2N-3]$.
@@
$=\binom{n}{N+1}\binom{n}{N-1}^{2}\binom{n}{k-N+1}\frac{4-(n-N-1)(2N-4-k)}{n^{2}}.$ (67)
@@
$=\binom{n}{N+1}\binom{n}{N}\binom{n}{N-1}\binom{n}{k-N}\frac{2-(n-N-1)(2N-2-k)}{n^{2}}.$ (68)
We now consider two subcases. If $k=2N-3$,  (67) is positive and the inequality $\tbinom{n}{N-1}\tbinom{n}{N-2}\geq\tbinom{n}{N}\tbinom{n}{N-3}$ ensures that the sum of (67) and (68) remains positive.
Otherwise, suppose that $k\in[N,2N-4]$. Again, minimizing the expression $\tbinom{n}{i}\tbinom{n}{j}$ under the constraints $i,j\leq N-2$ and $i+j=k$ requires $|i-j|$ to be largest possible (that is, $i=N-2$ and $j=k-N+2$ or vice versa). As a result, $\tbinom{n}{i}\tbinom{n}{j}\geq\binom{n}{N-2}\binom{n}{k-N+2}\geq\tbinom{n}{N-1}\binom{n}{k-N+1}\geq\tbinom{n}{N}\binom{n}{k-N}$ for all $i,j\leq N-2$ with $i+j=k$. Combining this observation with the inequality
@@
Case 3.3: $k=2N-2$.
Then, the coefficient in front of $x^{2N+k}$ in $P_{2}S_{1}-Q_{2}R_{1}$ is given by (68) and is positive.
Before we continue with the proof of J1J4, we simplify the expressions $P_{1}-Q_{1}=R_{2}-S_{2}$, $P_{2}-Q_{2}$ and $R_{1}-S_{1}$. First, using that, for every $\ell\geq 0$, $\binom{n-1}{\ell}=\tfrac{n-\ell}{n}\binom{n}{\ell}$, we obtain that
$P_{1}-Q_{1}=R_{2}-S_{2}=\sum_{i=0}^{N-1}\bigg{(}\binom{n}{N}\binom{n-1}{i}-\binom{n-1}{N}\binom{n}{i}\bigg{)}x^{N+i}=\sum_{i=0}^{N-1}\binom{n}{N}\binom{n}{i}\frac{N-i}{n}x^{N+i}.$ (70)
Furthermore, since $P_{2}$ (respectively, $Q_{2}$) is obtained from $P_{1}$ (respectively, $Q_{1}$) by replacing $N$ by $N+1$, we have
$P_{2}-Q_{2}=\sum_{i=0}^{N}\binom{n}{N+1}\binom{n}{i}\frac{N+1-i}{n}x^{N+i+1}.$ (71)
Simplifying $R_{1}-S_{1}$ follows a similar computation:
$R_{1}-S_{1}=\sum_{i=0}^{N-2}\bigg{(}\binom{n}{N-1}\binom{n-1}{i}-\binom{n-1}{N-1}\binom{n}{i}\bigg{)}x^{N+i-1}=\sum_{i=0}^{N-2}\binom{n}{N-1}\binom{n}{i}\frac{N-1-i}{n}x^{N+i-1}.$ (72)
Showing J1.
By I2, we already know that $P_{2}S_{2}\geq Q_{2}R_{2}$, so it remains to show that
$(P_{2}-Q_{2})B_{2}-(R_{2}-S_{2})A_{2}\geq 0.$ (73)
By (70) and (71), the desired inequality is equivalent to
$\sum_{i=0}^{N}\binom{n}{N+1}\binom{n}{i}\binom{n}{N}\binom{n-1}{N}\frac{N+1-i}{n}x^{3N+i+1}-\sum_{i=0}^{N-1}\binom{n}{N}\binom{n}{i}\binom{n}{N+1}\binom{n-1}{N+1}\frac{N-i}{n}x^{3N+i+2}\geq 0.$
By ignoring the summand for $i=0$ in the first sum above and changing the variable from $i\in[0,N-1]$ to $i-1\in[1,N]$ in the second sum, the last display is implied by the inequality
$\sum_{i=1}^{N}\bigg{(}\binom{n}{N+1}\binom{n}{i}\binom{n}{N}\binom{n-1}{N}-\binom{n}{N}\binom{n}{i-1}\binom{n}{N+1}\binom{n-1}{N+1}\bigg{)}\frac{N+1-i}{n}x^{3N+i+1}\geq 0,$
which is implied by the fact that, for every $i\in[N]$,
$\frac{\binom{n}{i}}{\binom{n}{i-1}}\cdot\frac{\binom{n-1}{N}}{\binom{n-1}{N+1}}=\frac{n-i+1}{i}\cdot\frac{N+1}{n-1-N}\geq\frac{n-N+1}{N}\cdot\frac{N+1}{n-1-N}>1.$
Showing J2.
By I2 and I3, we know respectively that $P_{2}S_{2}\geq Q_{2}R_{2}$ and $P_{1}S_{2}+P_{2}S_{1}\geq Q_{1}R_{2}+Q_{2}R_{1}$. Thus, in order to prove J3, it suffices to show that
$(P_{1}-Q_{1})B_{2}+(P_{2}-Q_{2})(B_{1}+B_{2})-(R_{1}-S_{1})A_{2}-(R_{2}-S_{2})(A_{1}+A_{2})\geq 0.$
Using (73) and recalling that $P_{1}-Q_{1}=R_{2}-S_{2}$, it is enough to show that
$(P_{1}-Q_{1})(B_{2}-A_{1})+(P_{2}-Q_{2})B_{1}-(R_{1}-S_{1})A_{2}\geq 0.$ (74)
To prove (74), observe that combining the fact that $A_{1}=2B_{2}$, (70) and a change of variable ensures that
$(P_{1}-Q_{1})(B_{2}-A_{1})=-(P_{1}-Q_{1})B_{2}$ $=-\sum_{i=0}^{N-1}\binom{n}{N}\binom{n}{i}\frac{N-i}{n}\binom{n}{N}\binom{n-1}{N}x^{3N+i}$
$=-\sum_{j=1}^{N}\binom{n}{N}\binom{n}{j-1}\frac{N+1-j}{n}\binom{n}{N}\binom{n-1}{N}x^{3N+j-1}.$ (75)
Furthermore, (71) and (72) ensure respectively that
$(P_{2}-Q_{2})B_{1}$ $=2\sum_{i=0}^{N}\binom{n}{N+1}\binom{n}{i}\frac{N+1-i}{n}\binom{n}{N-1}\binom{n-1}{N-1}x^{3N+i-1},$ (76)
$(R_{1}-S_{1})A_{2}$ $=\sum_{i=0}^{N-2}\binom{n}{N-1}\binom{n}{i}\frac{N-1-i}{n}\binom{n}{N+1}\binom{n-1}{N+1}x^{3N+i+1}$
$=\sum_{j=2}^{N}\binom{n}{N-1}\binom{n}{j-2}\frac{N+1-j}{n}\binom{n}{N+1}\binom{n-1}{N+1}x^{3N+j-1}.$ (77)
Henceforth, we adopt the convention that $\binom{b}{a}=0$ if $a<0$. Then, by combining (75), (76) and (77), the coefficient in front of $x^{3N+i-1}$ in (74) is non-negative if and only if
$2\binom{n}{N+1}\binom{n}{i}\binom{n}{N-1}\binom{n-1}{N-1}-\binom{n}{N-1}\binom{n}{i-2}\binom{n}{N+1}\binom{n-1}{N+1}-\binom{n}{N}\binom{n}{i-1}\binom{n}{N}\binom{n-1}{N}\geq 0$ (78)
For $i=0$, the inequality is obvious. For $i\in[N]$, observe that
$\frac{\binom{n}{i}}{\binom{n}{i-1}}\cdot\frac{\binom{n}{N-1}}{\binom{n}{N}}\cdot\frac{\binom{n}{N+1}\binom{n-1}{N-1}}{\binom{n}{N}\binom{n-1}{N}}=\frac{n-i+1}{i}\cdot\frac{N}{n-N+1}\cdot\frac{N}{N+1}\geq\frac{N}{N+1},$
and
$\frac{\binom{n}{N-1}}{\binom{n}{N}}\cdot\frac{\binom{n}{i-2}}{\binom{n}{i-1}}\cdot\frac{\binom{n}{N+1}}{\binom{n}{N}}\cdot\frac{\binom{n-1}{N+1}}{\binom{n-1}{N}}=\frac{N}{n-N+1}\cdot\frac{i-1}{n-i+2}\cdot\frac{n-N}{N+1}\cdot\frac{n-N-1}{N+1}\leq\frac{i-1}{N+1}\leq\frac{N-1}{N+1}.$
By combining the last two inequalities, it follows that the right hand side of (78) dominates
$\bigg{(}\frac{2N}{N+1}-\frac{N-1}{N+1}-1\bigg{)}\binom{n}{N}\binom{n}{i-1}\binom{n}{N}\binom{n-1}{N}=0,$
which finishes the proof of J2.
Showing J3.
By I1 and I3, we know respectively that $P_{1}S_{1}\geq Q_{1}R_{1}$ and $P_{1}S_{2}+P_{2}S_{1}\geq Q_{1}R_{2}+Q_{2}R_{1}$. Thus, in order to prove J3, it suffices to show that
$(P_{1}-Q_{1})(B_{1}+B_{2})+(P_{2}-Q_{2})B_{1}-(R_{1}-S_{1})(A_{1}+A_{2})-(R_{2}-S_{2})A_{1}\geq 0.$
By recalling that $P_{1}-Q_{1}=R_{2}-S_{2}$ and using (74), it is enough to show that
$(P_{1}-Q_{1})B_{1}-(R_{1}-S_{1})A_{1}\geq 0.$ (79)
@@
However, this inequality follows similarly to (73): indeed, note that $P_{2},Q_{2},R_{2},S_{2},A_{2},B_{2}$ are obtained respectively from $P_{1},Q_{1},R_{1},S_{1},\frac{1}{2}A_{1},\frac{1}{2}B_{1}$ by replacing $N$ with $N+1$.
Showing J4.
This inequality follows by combining I1 and (79). This concludes the proof of Lemma 2.6.